# Watch Your Language: Large Language Models and Content Moderation

### v1

@@ Abstract
Large language models (LLMs) have exploded in popularity due to their ability to perform a wide array of natural language tasks. Text-based content moderation is one LLM use case that has received recent enthusiasm, however, there is little research investigating how LLMs perform in content moderation settings. In this work, we evaluate a suite of modern, commercial LLMs (GPT-3, GPT-3.5, GPT-4) on two common content moderation tasks: rule-based community moderation and toxic content detection. For rule-based community moderation, we construct 95 LLM moderation-engines prompted with rules from 95 Reddit subcommunities and find that LLMs can be effective at rule-based moderation for many communities, achieving a median accuracy of 64% and a median precision of 83%. For toxicity detection, we find that LLMs significantly outperform existing commercially available toxicity classifiers. However, we also find that recent increases in model size add only marginal benefit to toxicity detection, suggesting a potential performance plateau for LLMs on toxicity detection tasks. We conclude by outlining avenues for future work in studying LLMs and content moderation.
@@ Introduction
Since the release of OpenAI’s ChatGPT in November 2022, publicly available large language models (LLMs) have exploded in popularity and achieved widespread recognition. Such models, like OpenAI’s GPT-suite and Google’s Bard, offer new and accessible ways for end-users to interface with machine learning and AI as well as provide developers with a new toolkit with which to build AI-powered technologies. Indeed, developers are already using LLMs for a variety of text-based tasks: Microsoft incorporated GPT-4 into Bing to enhance web search functionality, [^1] https://www.theverge.com/2023/3/15/23641683/microsoft-bing-ai-gpt-4-chatbot-available-no-waitlist Expedia released a GPT-3.5 powered chatbot to help end-users create travel plans, [^2] https://mashable.com/article/expedia-chatgpt-in-app-travel-booking and Slack incorporated a variety of LLMs to provide summaries of meetings and increase end-user productivity. [^3] https://slack.com/blog/news/introducing-slack-gpt
@@
One use case for LLMs that has achieved significant attention is automated content moderation. Although introducing AI to content moderation is not a new idea (Gillespie 2020), recent research has started investigating how best to incorporate LLMs into content moderation pipelines (Franco, Gaggi, and Palazzi 2023) and industry players alike are expressing enthusiasm for LLM-powered content moderation (Gilardi, Alizadeh, and Kubli 2023). OpenAI recently published an article outlining how GPT-4 could be used to build and evaluate content policies with test examples [^4] https://openai.com/blog/using-gpt-4-for-content-moderation to minimize time required to develop platform policy and reduce potential emotional burden for human reviewers. Given this recent attention, it behooves the research community to investigate the role that LLMs might play in automated content moderation, both in where LLMs excel and where there are opportunities for improvement.
@@
In this paper, we evaluate the effectiveness of several publicly available LLMs on two common content moderation tasks: rule-based community moderation and toxic content detection. For each task, we develop a test harness that combines real-world datasets with careful prompt design and evaluate each model against available ground truth. We focus on OpenAI’s GPT models (3, 3.5, 4) as exemplars of deployed language models. For the rule-based community moderation task, we build 95 LLM-powered moderation engines that are each prompted with community guidelines sourced from 95 subcommunities on Reddit. We prompt each LLM-engine to simulate moderation decisions on comments curated by Chandrasekharan et al. (2018) and compare how well LLM decisions align with human-moderator decisions. For toxic content detection, we construct a balanced dataset of 10K comments sourced by Kumar et al. (2021) and compare GPT-3, GPT-3.5, and GPT-4 performance with ground truth and other popular state-of-the-art commercially available baselines.
@@
We observe that LLMs can be effective at conducting rule-based moderation, achieving an accuracy of 64% and precision of 83% for the median subreddit. For some communities, GPT-3.5 achieves near human-moderator levels of performance: for example, the model achieves an accuracy of 82% and a precision of 95% on comments from r/movies. We investigate what rules GPT-3.5 evokes when simulating moderation decisions, and find that the model is more likely to moderate based on restrictive rules (e.g., those that restrict behavior) compared to prescriptive rules or formatting based rules, suggesting that framing policies as restrictions may be an effective strategy for moderators and policymakers seeking to use LLMs for moderation tasks. We also highlight potential pitfalls with using LLMs for moderation: we observe stark differences in performance when underlying models change over time, which can lead to unpredictable results.
@@
With toxicity detection, we observe that modern LLMs like GPT-3, GPT-3.5, and GPT-4 outperform existing commercial toxicity detection classifiers, with GPT-3.5 achieving the most balanced performance across all tested models (Accuracy of 0.74, F1 of 0.73). Interestingly, we find only marginal improvement in toxicity detection performance when comparing GPT-3, GPT-3.5, and GPT-4, suggesting that increased LLM-size may not always lead to increased benefits in a content moderation context. Still, LLMs are not perfect, and exhibit some significant issues: models often miss implicit toxicity, which has been a topic of recent study in toxicity detection (ElSherief et al. 2021), and, like existing classification-based approaches, at times cannot make accurate decisions without appropriate conversational context.
@@
Our results paint a tempered but optimistic view of introducing LLMs into content moderation contexts. LLMs provide some improvements over existing baselines but will still require human review and careful guardrails before widespread adoption. We conclude with a discussion of our results and chart a pathway for future research in this space. We will make our code and data available at publication time. We hope our results will serve useful to researchers and practitioners investigating how best to incorporate LLMs into content moderation pipelines.
@@
Google Jigsaw’s Perspective API (Google Jigsaw 2018) has emerged as the de facto classifer for detecting toxic content online, and it has been used in dozens of research papers to identify toxic content in isolation (Xia et al. 2020; Kumar et al. 2021; Lambert, Rajagopal, and Chandrasekharan 2022; Kumar et al. 2023). There has been significant additional research on improving detection beyond the Perspective API. Hanley et al., for example, demonstrate how a DeBERTa classifier trained with contrastive learning marginally outperforms the Perspective API (2023), and, in a forthcoming paper, He et al. show how prompting local LLMs (e.g., T5, GPT2-M and GPT2-L) can improve performance for a variety of toxicity related tasks (2024). While multiple studies have considered how older LLMs (i.e., GPT-3) can be used to identify hate speech (Chiu, Collins, and Alexander 2021; Schick, Udupa, and Schütze 2021), no work has studied modern LLMs (GPT-3.5, GPT-4) or compared how performance on toxicity detection has changed over time.
LLMs and Rule-Based Content Moderation
In this section, we study how LLMs can aid in rule-based moderation settings. We prompt OpenAI’s GPT-3.5 with lists of rules collected from 95 subreddits and ask the model to simulate moderation decisions for comments sourced from each subreddit. We evaluate the effectiveness of LLMs by comparing simulated decisions to ground-truth decisions made by human moderators.
@@ Methods
To compare how well LLMs align with human-provided moderation decisions, we leverage a dataset created by Chandrasekharan et al. (2018) that contains comments that were moderated from 100 popular subreddits between 2016–2017. Since their study was conducted, 5 subreddits have gone offline or been removed from the platform, so our study focuses on the remaining 95 subreddits. For each subreddit, we sample 500 comments that were removed by moderators as well as 500 comments that were not removed by moderators during the same time period, thus creating a balanced dataset of moderation decisions for each subreddit. These datasets serve as our baseline for rule-based evaluation.
@@
We construct a new zero-shot prompted LLM-engine for each subreddit we investigate. We begin by prompting the LLM with the description of the subreddit as well as the constructed rules string for each subcommunity. The full initial prompt text is:
@@
We then ask the model to evaluate whether it would moderate new unseen comments (yes/no), the rule text and rule numbers that the comment violates, and an explanation for the decision. Similarly to the toxic content detection task, we ask the model to return a JSON object for downstream analysis. The full prompt text is:
@@
Return a JSON object with four fields,
@@
which is a score from 1--5 on how
violative the comment is.
@@
We compare ground-truth moderation decisions with decisions made by each subreddit-specific tuned LLM on a number of binary performance metrics. For this experiment, we solely rely on OpenAI’s GPT3.5 for evaluation as it balances at-scale measurements (i.e., 100,000 queries in total) with cost. To understand why the LLM makes mistakes, we also perform a case study on mistakes created in the r/worldnews subreddit and provide a qualitative analysis of existing gaps in rule-based moderation.
@@
To begin, we measure the binary performance of each LLM-engine prompted with each subreddit’s rules compared to human moderators. GPT-3.5 achieves a median accuracy of 63.7%, a median precision of 83%, and a median recall of 39.8% across all subreddits in our dataset (Figure 1). For each metric, there is considerable spread: GPT-3.5 achieves high accuracy for subreddits like r/movies (81.5%) and r/OldSchoolCool (76.3%), with each of these also exhibiting high precision (94% and 82%, respectively). Such performance is not uniform across all subreddits. GPT-3.5 performs worse than a coin toss when simulating decisions from subreddits like r/askscience (Accuracy 38.5%) and r/AskHistorians (Accuracy 41.7%)—suggesting that LLMs are not always a useful community moderation-aid. Notably, both of these subreddits are centered around discussion with experts, and moderation decisions in these subcommunities often require significant context, which is a limitation of our evaluation setup (Gilbert 2020).
Figure 2: Threshold analysis for rule-based moderation—We show the effect of introducing LLM-created thresholds to rule-based moderation tasks across Accuracy, Precision, Recall, and F1. Thresholds in this context are not particularly sensitive, with thresholds from 1–3 achieving near identical results. As such, rating-based thresholding is likely a less effective strategy to enable more control over rule-based moderation than a simple binarized decision.
We next evaluate if thresholding (i.e., asking the model to provide a rating, and using that rating to binarize decisions) can be used to improve or tune rule-based performance, as has been useful in other content moderation contexts (Kumar et al. 2021). Figure 2 shows the binary metrics we use in this paper at each threshold from 0–4. Interestingly, we find that overall performance is not sensitive to thresholding—GPT-3.5 performed near identically for the middle three thresholds, with an accuracy of approximately 62.6% and an F1 of 0.52–0.54. At the highest threshold (t=4), precision greatly increased (0.94), but recall significantly diminished (0.06). Our results suggest that thresholding is not a useful strategy for rule-based moderation than in other contexts.
@@
Table 1: Longitudinal LLM performance—We show the median performance of GPT-3.5 across the first and second released models (from March 1, 2023 to June 13, 2023). We find that GPT-3.5 produces drastically different results, with new models significantly preferring precision over recall. Our results suggest that human-moderators may need to stay significantly in the loop as models grow and evolve.
@@
Given LLMs like GPT-3.5 are constantly undergoing changes due to model updates and deployment changes, we also sought to understand how stable model performance is at rule-based moderation tasks over time. Table 1 shows the median performance across all subreddits in our dataset from the first released GPT-3.5 model (released on March 1, 2023) and the second released model (released on June 13, 2023). Performance significantly changes between the two models, despite there being no other differences in prompts or analyzed comments. F1 performance decreased from 0.68 to 0.53, primarily due to a decrease in median recall: which similarly decreased from 0.63 to 0.40. Median precision, however, increased from 0.75 to 0.83—suggesting that the newer model favored reducing false positives over reducing false negatives. Although it is unclear why the underlying GPT-3.5 model shifted in this direction, our results highlight that content moderators wishing to incorporate LLMs into their moderation workflows will need to consider model stability as an additional factor when deploying these tools.
@@
LLM decisions are not only governed by the content, but also by the types of rules that are available when making moderation decisions. As part of our prompt design, we ask the LLM to express which rules are violated when it flags a comment for moderation. We use these rules to study what kinds of rules LLMs may be better at expressing over other rules, which can inform how LLMs can be better used in practice by community moderators. To increase our descriptive power for each rule, we replicate the work of Fiesler et al. (2018), who created a model and taxonomy for classifying Reddit rules into several broad categories. We classify each subreddit rule into three categories: restrictive rules (i.e., those that restrict types of behavior), prescriptive rules (i.e., those that tell someone what to do), and format related rules. We detail our model replication process in the appendix.
GPT-3.5 expresses a significant fraction of rules: of the 562 rules in our dataset, GPT-3.5 expressed 467 (83.1%) when making moderation decisions. The median subreddit has 88% of its rules exercised by the LLM, with 49% of subreddits having 100% of their rules expressed by the LLM. Still, there is significant distributional spread, with some subreddits having only a small fraction of their rules expressed. For example, r/AskTrumpSupporters, a subreddit for asking questions to those that supported Donald Trump’s presidency, only has 3 of its 7 rules expressed.
@@
% Comments
Mistaken Hate
47%
Correct
23%
Context
20%
@@
Table 3: Explaining errors in moderating r/worldnews—We show the most frequent explanations for errors generated by GPT-3.5 when moderating r/worldnews. False positives are largely due to incorrectly flagging a comment as hate speech, especially when identity language appears (47%). False negatives occur most often when there is missing additional context that needs to be included to make a moderation decision (54%).
We complement our quantitative analyses with a qualitative analysis on errant decisions made by GPT-3.5 on the subreddit r/worldnews, which is a subreddit dedicated to major news around the world. We chose this particular subreddit as it is popular (33M subscribers), it well-balances false positives (33%) and false negatives (67%) to analyze, and tends to more general discussion versus niche subcommunity interests. r/worldnews has only 3 rules that apply to comments, two of which are related to personal attacks and hate speech, and one which is related to meme-content. We manually investigate the 244 comments that produced either a false negative or false positive, iteratively develop a codebook for explaining each error, and manually code each comment with an explanation. Table 3 shows a distribution of common explanations for false positives and false negatives generated by model. We discuss each in detail below:
@@
The primary reason for false positives is due to GPT-3.5 mistakenly identifying a comment as hate speech (47%). This often occurs when identity language is prevalent in the comment and the sentiment is not overtly clear. For example, the comment “Are you aware that there are other businessmen in the world besides your Jew Boogeyman?” was flagged as violative of hate speech rules, despite the fact that it is attempting to rebuff the original commenter and defend Jewish people from unwanted stereotypes. Given that hate speech and hateful language is part of the Reddiquette, [^6] https://support.reddithelp.com/hc/en-us/articles/205926439-Reddiquette which governs platform-wide norms and behaviors, errors produced by LLMs like these will have a wide impact on all communities wishing to curb and moderate hate speech. Interestingly, we also observe 23% of “false positives” where we agree with the model decision over the human-rater. For example, the comment “Lol you’re a fucking idiot” was not flagged by human moderators, but clearly violates the second rule of r/worldnews. The existence of discrepancies like this is encouraging, as it suggests LLMs can also potentially be used as a double-check aid for human moderators who may also make mistakes when moderating content.

### v2

# Watch Your Language: Investigating Content Moderation with Large Language Models

@@ Abstract
Large language models (LLMs) have exploded in popularity due to their ability to perform a wide array of natural language tasks. Text-based content moderation is one LLM use case that has received recent enthusiasm, however, there is little research investigating how LLMs perform in content moderation settings. In this work, we evaluate a suite of commodity LLMs on two common content moderation tasks: rule-based community moderation and toxic content detection. For rule-based community moderation, we instantiate 95 subcommunity specific LLMs by prompting GPT-3.5 with rules from 95 Reddit subcommunities. We find that GPT-3.5 is effective at rule-based moderation for many communities, achieving a median accuracy of 64% and a median precision of 83%. For toxicity detection, we evaluate a suite of commodity LLMs (GPT-3, GPT-3.5, GPT-4, Gemini Pro, LLAMA 2) and show that LLMs significantly outperform currently widespread toxicity classifiers. However, recent increases in model size add only marginal benefit to toxicity detection, suggesting a potential performance plateau for LLMs on toxicity detection tasks. We conclude by outlining avenues for future work in studying LLMs and content moderation.
@@ Introduction
Since the release of OpenAI’s ChatGPT in November 2022, publicly available large language models (LLMs) have exploded in popularity and achieved widespread recognition. Such systems, like OpenAI’s GPT-suite and Google’s Gemini, offer new and accessible ways for end-users to interface with AI as well as provide developers with a new toolkit with which to build AI-powered technologies. Indeed, developers are already using LLMs for a variety of text-based tasks: Microsoft incorporated GPT-4 into Bing to enhance web search functionality, [^1] https://www.theverge.com/2023/3/15/23641683/microsoft-bing-ai-gpt-4-chatbot-available-no-waitlist Expedia released a GPT-3.5 powered chatbot to help end-users create travel plans, [^2] https://mashable.com/article/expedia-chatgpt-in-app-travel-booking and Slack incorporated a variety of LLMs to provide summaries of meetings and increase end-user productivity. [^3] https://slack.com/blog/news/introducing-slack-gpt
@@
One use case for LLMs that has achieved significant attention is automated content moderation. Although using AI for content moderation is not a new idea (Gillespie 2020), recent research has begun investigating how best to incorporate LLMs into moderation pipelines (Franco, Gaggi, and Palazzi 2023) and industry players alike are expressing enthusiasm for LLM-powered moderation (Gilardi, Alizadeh, and Kubli 2023). For example, OpenAI recently published an article outlining how GPT-4 could be used to build and evaluate content policies with test examples [^4] https://openai.com/blog/using-gpt-4-for-content-moderation to minimize time required to develop platform policy and reduce emotional burden for human reviewers. Given this recent attention, it behooves the research community to study the role that LLMs might play in automated content moderation, both in where LLMs excel and where there are opportunities for improvement.
@@
In this paper, we evaluate the effectiveness of several publicly available LLMs on two common content moderation tasks: rule-based community moderation and toxic content detection. For each task, we develop a test harness that combines real-world datasets with careful prompt design and evaluate each model against available ground truth. We focus on OpenAI’s GPT models (3, 3.5, 4), Google’s Gemini Pro, and Meta’s LLAMA as recent exemplars of LLMs deployed in practice for our evaluation. For the rule-based community moderation task, we instantiate 95 subcommunity specific LLMs by prompting GPT-3.5 with community guidelines sourced form 95 subcommunities on Reddit. We then prompt each subreddit-specific model to simulate moderation decisions on real comments sourced from each community and compare how well LLM decisions align with human-moderator decisions. For toxic content detection, we construct a balanced dataset of 10K comments sourced by Kumar et al. (2021) and compare GPT-3, GPT-3.5, GPT-4, Gemini Pro, and LLAMA 2 performance with ground truth and other popular state-of-the-art commercially available baselines.
@@
We observe that GPT-3.5 is effective at conducting rule-based moderation, achieving an accuracy of 64% and precision of 83% for the median subreddit. For some communities, GPT-3.5 achieves near human-moderator levels of performance: for example, the model achieves an accuracy of 82% and a precision of 95% on comments from r/movies. We investigate what rules GPT-3.5 evokes when simulating moderation decisions, and find that the model is more likely to moderate based on restrictive rules (e.g., those that restrict behavior) compared to prescriptive rules or formatting based rules, suggesting that framing policies as restrictions may be an effective strategy for moderators and policymakers seeking to use LLMs for moderation tasks. We also highlight potential pitfalls with using LLMs for moderation: we observe stark differences in performance when underlying models change over time, which can lead to unpredictable results.
@@
With toxicity detection, we observe that modern LLMs outperform existing commercial toxicity detection classifiers, with GPT-3.5 achieving the most balanced performance across all tested models (Accuracy of 0.73, F1 of 0.75). Interestingly, we find only marginal improvement in toxicity detection performance when comparing GPT-3, GPT-3.5, GPT-4, and Gemini Pro, suggesting that increased LLM-size may not always lead to increased benefits in a content moderation context. Still, LLMs are not perfect, and exhibit some significant issues: models often miss implicit toxicity, which has been a topic of recent study in toxicity detection (ElSherief et al. 2021), and, like existing classification-based approaches, at times cannot make accurate decisions without appropriate world context.
@@
We view our results as a tempered by optimistic first step in introducing LLMs into content moderation contexts. LLMs provide some improvements over existing baselines but will still require human review and careful guardrails before widespread adoption. We conclude with a discussion of our results and chart a pathway for future research in this space. We will make our code and data available at publication time; data is available anonymously with this submission. [^5] https://anonymous.4open.science/r/llm-content-mod-1F4B We hope our results will serve useful to researchers and practitioners investigating how best to incorporate LLMs into content moderation pipelines.
@@
Google Jigsaw’s Perspective API (Google Jigsaw 2018) has emerged as the de facto classifer for detecting toxic content online, and it has been used in dozens of research papers to identify toxic content in isolation (Xia et al. 2020; Kumar et al. 2021; Lambert, Rajagopal, and Chandrasekharan 2022; Kumar et al. 2023). There has been significant additional research on improving detection beyond the Perspective API. Hanley et al., for example, demonstrate how a DeBERTa classifier trained with contrastive learning marginally outperforms the Perspective API (2023), and, in a forthcoming paper, He et al. show how prompting local LLMs (e.g., T5, GPT2-M and GPT2-L) can improve performance for a variety of toxicity related tasks (2024). Multiple studies have considered how older LLMs (i.e., GPT-3) can be used to identify hate speech (Chiu, Collins, and Alexander 2021; Schick, Udupa, and Schütze 2021), and recent research has studied how effective GPT-3.5 is at detecting hate speech in English (Li et al. 2023) and Portugese (Oliveira et al. 2023).
Rule-Based Content Moderation
In this section, we study how LLMs can aid in rule-based moderation settings by instantiating 95 subreddit-specific LLMs with lists of community guidelines sourced from 95 subreddits. We ask each subreddit-specific model to simulate moderation decisions for comments sourced from each subreddit and evaluate their performance.
@@ Methods
To collect ground-truth moderation decisions, we leverage a dataset created by Chandrasekharan et al. (2018) that contains comments that were moderated from 100 popular subreddits between 2016–2017. Since their study was conducted, 5 subreddits have gone offline or been removed from the platform, so our study focuses on the remaining 95 subreddits. For each subreddit, we sample 500 comments that were removed by moderators as well as 500 comments that were not removed by moderators during the same time period, thus creating a balanced dataset of moderation decisions for each subreddit. These datasets serve as our baseline for rule-based evaluation.
@@
We instantiate a new zero-shot prompted LLM for each subreddit we investigate. We begin by prompting the LLM with the description of the subreddit as well as the constructed rules string for each subcommunity. The full initial prompt text is:
@@
We then ask the model to evaluate whether it would moderate new unseen comments (yes/no), the rule text and rule numbers that the comment violates, and an explanation for the decision. We ask the model to return a JSON object with consistent indexing for ease of downstream analysis. The full prompt text is:
@@
Return a JSON object with five fields,
@@
which is a score from 1-5 on how
violative the comment is, and "explanation"
which provides a reason for your decision.
@@
We compare ground-truth moderation decisions with decisions made by each subreddit-specific LLM on a number of binary performance metrics. For this experiment, we solely evaluate OpenAI’s GPT-3.5 as an exemplar LLM as it balances at-scale measurements (i.e., 100,000 queries in total) with cost. We set the “temperature” to 0 to ensure our results are deterministic across repeated queries. We also conduct a case study comparison between with Google’s Gemini Pro for a small number of subreddits to examine if our results generalize to other LLMs. To understand why each subreddit-specific model makes mistakes, we also perform a case study on mistakes created in the r/worldnews subreddit and provide a qualitative analysis of existing gaps in rule-based moderation.
@@
To begin, we measure the binary performance of each subreddit-specific LLM when compared to human moderators. GPT-3.5 achieves a median accuracy of 63.7%, a median precision of 83%, and a median recall of 39.8% across all subreddits in our dataset (Figure 1). For each metric, there is considerable spread: GPT-3.5 achieves high accuracy for subreddits like r/movies (81.5%) and r/OldSchoolCool (76.3%), with each of these also exhibiting high precision (0.95 and 0.82, respectively). Such performance is not uniform across all subreddits. GPT-3.5 performs worse than a coin toss when simulating decisions from subreddits like r/askscience (Accuracy 38.5%) and r/AskHistorians (Accuracy 41.7%)—suggesting that LLMs are not always a useful community moderation-aid. Notably, both of these subreddits are centered around discussion with experts, and moderation decisions in these subcommunities often require significant world context, which is a limitation of our evaluation setup (Gilbert 2020).
Prompt Sensitivity
Subreddit
Metrics Base Platform Rules CoT
r/movies
Accuracy 0.81 0.82 0.81
Precision 0.95 0.92 0.92
Recall 0.68 0.72 0.7
F1 0.79 0.81 0.8
r/IAmA
Accuracy 0.64 0.64 0.62
Precision 0.84 0.82 0.78
Recall 0.37 0.4 0.38
F1 0.52 0.53 0.51
r/askscience
Accuracy 0.38 0.39 0.4
Precision 0.87 0.95 0.95
Recall 0.11 0.11 0.11
F1 0.19 0.19 0.19
Table 1: Rule-based moderation prompt sensitivity—Rule-based content moderation is robust to differing prompt strategies; overall performance is comparable across the three strategies we tested for each subreddit.
LLM output is often highly sensitive to input prompts (White et al. 2023; OpenAI 2023). As such, we also examine how our results vary with different prompting strategies. We evaluate two prompt strategies in addition to our base prompt: providing more world-context by embedding platform rules, and chain-of-thought prompting, which is known to elicit reasoning and often better performance from commodity LLMs (Wei et al. 2022). For the embedded world-context prompt, we prepend to the community rules a list of comment-related rules from the Reddiquette [^7] https://support.reddithelp.com/hc/en-us/articles/205926439-Reddiquette, which is a list of platform-wide rules that most subreddits adhere to. For chain-of-thought prompting, we include the phrase “Let’s think step by step.” at the end of our prompt. For this experiment, we evaluate comments from three subreddits: r/movies, r/IAmA, and r/askscience; these represent communities where GPT-3.5 performed the best, median, and worst in our dataset. We select a subsample of subreddits as the cost of running every subreddit query with every prompt variant was prohibitively expensive (a single run costs upwards of \$200 US dollars).
In general, our performance results are robust to different prompt strategies (Table 1). The inclusion of the Reddiquette and CoT-prompting does marginally shift the tradeoff made between precision and recall: in general, embedding the Reddiquette or using CoT prompting resulted in slightly lower precision and slightly higher recall when compared to the base prompt, however, aggregate performance is near identical between the three. These observations are flipped when moderating r/askscience—including platform rules and CoT prompting increased precision meaningfully (from 0.87 to 0.95) without any change to recall, suggesting that for some communities, adding context to prompts can provide significant utility.
@@
Table 2: Longitudinal LLM performance—We show the median performance of GPT-3.5 across the first and second released models (from March 1, 2023 to June 13, 2023). We find that GPT-3.5 produces drastically different results, with new models significantly preferring precision over recall. Our results suggest that human moderators need to stay significantly in the loop as models grow and evolve.
@@
Given systems like GPT-3.5 are constantly undergoing changes due to model updates and deployment changes, we also sought to understand how stable model performance is at rule-based moderation tasks over time. Table 2 shows the median performance across all subreddits in our dataset from the first released GPT-3.5 model (released on March 1, 2023) and the second released model, which is the one we report on in this paper (released on June 13, 2023). Performance significantly changes between the two models, despite there being no other differences in prompts or analyzed comments. F1 performance decreased from 0.68 to 0.53, primarily due to a decrease in median recall: which similarly decreased from 0.63 to 0.40. Median precision, however, increased from 0.75 to 0.83—suggesting that the newer model favored reducing false positives over reducing false negatives. Although it is unclear why the underlying GPT-3.5 model shifted in this direction, our results highlight that content moderators wishing to incorporate LLMs into their moderation workflows will need to consider model stability as an additional factor when deploying these tools.
@@
LLM decisions are not only governed by the content, but also by the types of rules that are available when making moderation decisions. As part of our prompt design, we ask the LLM to express which rules are violated when it flags a comment for moderation. We use these rules to study what kinds of rules LLMs may be better at expressing over other rules, which can inform how LLMs can be better used in practice by community moderators. To increase our descriptive power for each rule, we replicate the work of Fiesler et al. (2018), who created a model and taxonomy for classifying Reddit rules into broad categories. We classify each subreddit rule into three categories: restrictive rules (i.e., those that restrict types of behavior), prescriptive rules (i.e., those that tell someone what to do), and format related rules. We detail our model replication process in the Appendix.
GPT-3.5 expresses a significant fraction of rules: of the 562 rules in our dataset, GPT-3.5 expressed 467 (83.1%) when making moderation decisions. The median subreddit has 88% of its rules expressed by GPT-3.5, with 49% of subreddits having 100% of their rules expressed by the LLM. Still, there is significant distributional spread, with some subreddits having only a small fraction of their rules expressed. For example, r/AskTrumpSupporters, a subreddit for asking questions to those that supported Donald Trump’s presidency, only has 3 of its 7 rules expressed.
@@
Moderating Subreddits in Realtime
Our analysis thus far has focused on evaluating LLMs in a balanced class scenario to provide a benchmark for performance. To understand how LLMs perform in a more realistic scenario, we applied our subreddit-specific LLMs to realtime decisions made by community moderators. To capture this, we collected our own realtime dataset by repeating the procedure in Chandrasekharan et. al (Chandrasekharan et al. 2018). We first stream comments from the 95 subreddits of interest, and second, query which of those comments were eventually removed (denoted with the text of the comment changing to “[removed]”) by moderators after 24-hours. We collected realtime moderation decisions for a 3 week period between November 11 2023 and and December 18 2023.
Given the cost constraints with querying GPT-3.5, we did not run every comment through each subreddit-specific LLM. Rather, we sampled moderated and unmoderated comments from each subreddit at each subreddit’s moderation rate. In general, moderation rates were low, with an average of 2% of comments moderated and rates ranging from as small as 0.2% (r/CFB) to at most 15.7% (r/legaladvice). We selected a sample size for each subreddit large enough to ensure our results are robust at an error rate of 5% and a CI of 95%, and did not analyze any subreddits for which we were not able to achieve sufficient comment volume. Ultimately, we curated and analyzed realtime moderation decisions from 62 of our 95 subreddits.
The realtime moderation task is more challenging for each subreddit-specific LLM than a balanced task, with the average ROC_AUC decreasing from 0.67 to 0.6 across all comments. Some subreddits, however, saw greater performance in a realtime setting: for example, the ROC_AUC for r/gonewild increased from 0.56 to 0.61. Accuracy for the realtime task increases from 70% to 90%, however, this is largely due to a significant increase in unmoderated comments with significant class imbalance. Average recall drops from 43% to 29% in the realtime setting. Precision also falls drastically from 83% to just 6.2%—however, prior work has well documented the noisiness of realtime moderation decisions from Reddit communities. For example, Chandrasekharan et al. found that for Crossmod, a realtime moderation system they designed, Reddit moderators would have moderated 95% of unmoderated comments Crossmod flagged had the tool existed (Chandrasekharan et al. 2019), emphasizing that precision is a poor metric for evaluating realtime moderation tools on Reddit. Our results echo this idea, and highlight the need to evaluate systems in-context (i.e., in collaboration with moderators) to understand the practical impact such systems might offer.
@@
Model Correct
48%
Mistaken Attack
17%
Generalization
14%
Hate Speech
8%
@@
Table 4: Explaining errors in moderating r/worldnews—We show the most frequent explanations for errors generated by GPT-3.5 when moderating r/worldnews. We agree with model decisions for 48% of false positives and 16% of false negatives. In general, models often make mistakes due to missing context.
We complement our quantitative analyses with a qualitative analysis on errant decisions made by GPT-3.5 on the subreddit r/worldnews, which is a subreddit dedicated to major news around the world. We chose this particular subreddit as it is popular (33M subscribers), it well-balances false positives (33%) and false negatives (67%) to analyze, and tends to more general discussion versus niche subcommunity interests. r/worldnews has only 3 rules that apply to comments, two of which are related to personal attacks and hate speech, and one which is related to meme-content. Two independent raters manually investigated the 244 comments that produced either a false negative or false positive, iteratively developed a codebook for explaining each error, and coded each comment. Inter-rater agreement was high (Cohen’s Kappa 0.89), and disagreements were manually resolved. Due to noisiness in subreddit moderation, we first determined whether Table 4 shows a distribution of common explanations for false positives and false negatives produced by the model. We discuss each in detail below:
@@
We observe that for 48% of false positives, we agree with the model decision over the human rater. For example, the comment “Lol you’re a fucking idiot” was not flagged by human moderators, but clearly violates the second rule of r/worldnews and the Reddiquette. The existence of discrepancies like this is encouraging as it suggests LLMs can potentially be used as a double-check aid for human moderators who may inadvertently make mistakes. The next top reasons are when GPT-3.5 mistakenly identifies a comment as an attack (17%), as a generalization (14%), or as hate speech (8%). For example, the comment “Are you aware that there are other businessmen in the world besides your Jew Boogeyman?” was flagged as violative of hate speech rules, despite the fact that it is attempting to rebuff the original commenter and defend Jewish people from unwanted stereotypes. Given that hate speech and hateful language is part of the Reddiquette, errors produced by LLMs like these will have a wide impact on all communities wishing to curb and moderate hate speech.